# Suppressing spurious transitions using spectrally balanced pulse

### v1

@@ Acknowledgements.
The authors would like to thank Yasunobu Nakamura and Peng Zhao for fruitful discussions. This work was supported by the National Natural Science Foundation of China (Grants No. 12404558, No. 12322413, No. 92365206, and No. 92476206), Innovation Program for Quantum Science and Technology (Grants No.2021ZD0301802).
Supplementary Materials for “Suppressing spurious transitions using spectrally balanced pulse”
Supplementary Materials for “Suppressing spurious transitions using spectrally balanced pulse”
@@ III.2 General flow for the calibration of $\rm\sqrt{X}$ gate
[…]
@@
3. Prepare the initial state in $\left|\rm g\right\rangle$, apply $24n+$ 2 gates to be calibrated, and scan the driving amplitude. Take the value of the driving amplitude that corresponds to the maximum of $P^{\rm e}$ as the updated $\Omega$.
@@ III.3 Calibration of $\rm\sqrt{X}$ gate with dual-DRAG
Figure 5: Calibration processes with the pulse train method. Measured $P^{\rm e}$ of $\rm Q_{1}$ using the same detection circuit as in Fig. 3(b) of the main text, but with dual-DRAG, plotted against the waiting time $\tau$ and the spectral hole-related detuning $\Delta$. The white dashed line indicates the optimized $\Delta$ value used for subsequent experiments.
@@ IV.1 ZX crosstalk
where $A^{\prime}=\frac{\Omega^{2}g}{-2\Delta_{0}(\Delta_{0}^{2}-\Omega^{2})}$, $B^{\prime}=\frac{\Omega g}{-2(\Delta_{0}^{2}-\Omega^{2})}$, $\theta=\Delta_{0}(T_{g}+T_{b})$. The results show that, when $\theta=m\pi$ ($m\in N$), the term of $|\frac{\sin{n\theta}}{\sin{\theta}}|$ is in the peak value. However, when $\theta=2k\pi$ ($k\in N$), there is $01-e^{-i\theta}=$ 0, then only when $\theta=(2k+1)\pi$, we can detect the peaks in the states $\left|\rm ge\right\rangle$ and $\left|\rm ee\right\rangle$. The buffer time between two adjacent peaks is $\Delta T_{b}=|2\pi/\Delta_{0}|$. We can also know that, for our sequence, when the initial state is $\left|\rm gg\right\rangle$, the peaks of $\left|\rm ee\right\rangle$ represents the interaction of the first-order process between $\left|\rm gg\right\rangle$ and $\left|\rm ge\right\rangle$ or $\left|\rm eg\right\rangle$ and $\left|\rm ee\right\rangle$, while the peaks of $\left|\rm ge\right\rangle$ indicates the second-order interaction between $\left|\rm gg\right\rangle$ and $\left|\rm ee\right\rangle$ or $\left|\rm ge\right\rangle$ and $\left|\rm eg\right\rangle$.
@@ V Analysis for DRAG corrections
According to the experimental and numerical simulation results, the maximum amplitude of the driving pulse is about 30 MHz, then we assume the maximum value of $\Omega^{(0)}/\Delta_{0}\approx 3/4$, the detuning is about $-1.39[\Omega^{(0)}]^{2}/\Delta_{0}$. The item related to the first-order quantum crosstalk is $H^{V}[1,2]=-0.89[\Omega^{(0)}]^{3}g/2\Delta_{0}^{3}$ and $H^{V}[4,3]=-H^{V}[1,2]$, compared with the original item in $H_{r}[1,2]$, there is $H^{V}[1,2]/H_{r}[1,2]=0.89[\Omega^{(0)}]^{2}/\Delta_{0}^{2}\approx 0.5$, which is not far less than 1. That is the reason for the loss of efficacy of the traditional DRAG correction. There is one thing to note, even without considering the phase correction and setting $\eta=0$, the value of $H^{V}[1,2]/H_{r}[1,2]$ about $0.28$, the suppression for the crosstalk induced transitions is still limited.
@@ VI Fitting function for the excitation rate
[…]
@@
where $p_{1}(m-1)$ is the population on the excited state after $m-1$ Clifford gates. We can get the population of the excited state in $\rho_{m}$ as
@@ VII Explanation of the inter-pulse interference effects
[…]
@@ VIII Comparison of different DRAG sets
To compare the suppression of spurious transitions with different DRAG sets, we present the experimental results for RB and pulse train with four different DRAG sets, as shown in Fig. 8. The ExPC is significantly reduced with the recursive DRAG scheme, and the best suppression occurs with the optimized dual-DRAG, using the DRAG set $\{\alpha,\Delta_{\rm opt},-\Delta_{\rm opt}\}$, as shown in Fig. 8(b). The significant suppression is further confirmed in Fig. 8(c). The irregular lines in the top panel may be attributed to noise in the system.
@@ X Suppression of the spurious transitions
XI Leakage suppression with recursive DRAG

### v2

@@ Acknowledgements.
The authors would like to thank Yu He, Yasunobu Nakamura and Peng Zhao for fruitful discussions. This work was supported by Beijing Natural Science Foundation (Grants No. JQ25014), National Natural Science Foundation of China (Grants No. 12404558, No. 12322413, No. 92365206, and No. 92476206), Innovation Program for Quantum Science and Technology (Grants No.2021ZD0301802).
Data availability — The data that support the findings of this Letter are available from the authors upon reasonable request.
Supplemental Materials for “Suppressing spurious transitions using spectrally balanced pulse”
@@ III.2 General flow for the calibration of $\rm\sqrt{X}$ gate
1. Prepare the qubit in the initial state $\left|\text{g}\right\rangle$, apply the gate to be calibrated, and scan the driving power. The power value corresponding to the final state being $\frac{1}{\sqrt{2}}(\left|\rm g\right\rangle+\left|\rm e\right\rangle)$ is taken as the initial driving amplitude $\Omega_{0}$.
2. Prepare the qubit in the initial state $\left|\text{g}\right\rangle$, and implement two gates to be calibrated, with a VZ compensation in between, having an angle $\phi$. Scan the value of $\phi$, and take the value that corresponds to the maximum probability of the final state being in $\left|\rm e\right\rangle$ as the initial angle $\phi_{0}$ for the VZ compensation.
@@
3. Prepare the initial state in $\left|\rm g\right\rangle$, apply $4n+2$ gates to be calibrated, and scan the driving amplitude. Take the value of the driving amplitude that corresponds to the maximum of $P^{\rm e}$ as the updated $\Omega$.
@@ III.3 Calibration of $\rm\sqrt{X}$ gate with dual-DRAG
Figure 5: Calibration processes with the interference error filter method. Measured $P^{\rm e}$ of $\rm Q_{1}$ using the same detection circuit as in Fig. 3(b) of the main text, but with dual-DRAG, plotted against the waiting time $\tau$ and the spectral hole-related detuning $\Delta$. The white dashed line indicates the optimized $\Delta$ value used for subsequent experiments.
@@ IV.1 ZX crosstalk
where $A^{\prime}=\frac{\Omega^{2}g}{-2\Delta_{0}(\Delta_{0}^{2}-\Omega^{2})}$, $B^{\prime}=\frac{\Omega g}{-2(\Delta_{0}^{2}-\Omega^{2})}$, $\theta=\Delta_{0}(T_{g}+T_{b})$. The results show that, when $\theta=m\pi$ ($m\in N$), the term of $|\frac{\sin{n\theta}}{\sin{\theta}}|$ is in the peak value. However, when $\theta=2k\pi$ ($k\in N$), there is $1-e^{-i\theta}=0$, then only when $\theta=(2k+1)\pi$, we can detect the peaks in the states $\left|\rm ge\right\rangle$ and $\left|\rm ee\right\rangle$. The buffer time between two adjacent peaks is $\Delta T_{b}=|2\pi/\Delta_{0}|$. We can also know that, for our sequence, when the initial state is $\left|\rm gg\right\rangle$, the peaks of $\left|\rm ee\right\rangle$ represents the interaction of the first-order process between $\left|\rm gg\right\rangle$ and $\left|\rm ge\right\rangle$ or $\left|\rm eg\right\rangle$ and $\left|\rm ee\right\rangle$, while the peaks of $\left|\rm ge\right\rangle$ indicates the second-order interaction between $\left|\rm gg\right\rangle$ and $\left|\rm ee\right\rangle$ or $\left|\rm ge\right\rangle$ and $\left|\rm eg\right\rangle$.
@@ V Analysis for DRAG corrections
According to the experimental and numerical simulation results, the maximum amplitude of the driving pulse is about 30 MHz, then we assume the maximum value of $\Omega^{(0)}/\Delta_{0}\approx 3/4$, the detuning is about $1.39[\Omega^{(0)}]^{2}/\Delta_{0}$. The item related to the first-order quantum crosstalk is $H^{V}[1,2]=-0.89[\Omega^{(0)}]^{3}g/2\Delta_{0}^{3}$ and $H^{V}[4,3]=-H^{V}[1,2]$, compared with the original item in $H_{r}[1,2]$, there is $H^{V}[1,2]/H_{r}[1,2]=0.89[\Omega^{(0)}]^{2}/\Delta_{0}^{2}\approx 0.5$, which is not far less than 1. That is the reason for the loss of efficacy of the traditional DRAG correction. There is one thing to note, even without considering the phase correction and setting $\eta=0$, the value of $H^{V}[1,2]/H_{r}[1,2]$ about $0.28$, the suppression for the crosstalk induced transitions is still limited.
@@
In reference [17], based on waveform analysis, the Fourier Ansatz Spectrum Tuning Derivative Removal by Adiabatic Gate (FAST DRAG) and higher-derivative (HD) DRAG methods have been developed to reduce leakage to the second excited state. This is achieved with a gate duration of 6.25 ns and a leakage error of $3\times 10^{-5}$. In their work, they considered a special case of the HD DRAG method with in-phase and quadrature envelopes, as described by:
$\Omega_{I}(t)$ $=$ $A[g(t)+\beta_{2}\ddot{g}(t)]$ (40)
$\Omega_{Q}(t)$ $=$ $-\frac{A\beta}{\alpha}[\dot{g}(t)+\beta_{2}\dddot{g}(t)]$ (41)
The in-phase term in Eq.(40) is similar to our Eq.(35). The effectiveness of the in-phase envelope in HD DRAG, can be explained by Eq. (35). The second derivative term corresponds precisely to the two symmetric DRAG terms, and its coefficient determines the position of the blocking frequency in the spectrum. Therefore, the pulse for $\Omega_{I}$ is equivalent to an initial pulse of $Ag(t)$, plus a pair of spectrally symmetric DRAG terms, one of which resides at the frequency $f_{\rm ef}$. Meanwhile, $\Omega_{Q}$ represents the third DRAG, which also resides at the frequency $f_{\rm ef}$. This explains why HD DRAG outperforms conventional DRAG in suppressing leakage.
To provide a more detailed explanation of the frequency domain characteristics of different waveforms, we present the frequency domain of a 6-ns pulse with a frequency of 4.2 GHz and an original pulse envelope of $\Omega^{(0)}=\sin^{4}{\left(\frac{\pi t}{t_{g}}\right)}$ in Fig. 7. The blue line represents the dual-DRAG pulse in the frequency domain, with the DRAG detuning occurring at the frequencies of $\pm\alpha$, where $\alpha/2\pi=-210\,\text{MHz}$ is the typical frequency of the anharmonicity. The green line shows the dual-DRAG pulse with the addition of a third DRAG at $\alpha$, which is similar to the third-derivative HD DRAG method discussed in reference [17]. The red line represents a pulse with two DRAGs at the same DRAG detuning of $\alpha$. In this case, the drive frequency is shifted by about 170 MHz, whereas the frequency shift of the pulse shown in green is only about 50 MHz. A smaller frequency shift simplifies the calibration process and improves quantum gate fidelity. The advantage of the pulse with a smaller frequency shift arises from the symmetric DRAG at $\alpha$, which is the key point discussed in our paper.
Figure 7: Frequency domain representation of the pulse with three different DRAG settings. The central frequency of the initial pulse without DRAG is $\omega_{0}/2\pi=4.2~\mathrm{GHz}$, with a pulse duration of $t_{g}=6~\mathrm{ns}$ and anharmonicity $\alpha/2\pi=-210~\mathrm{MHz}$. The blue line shows the pulse with dual-DRAG (DRAG set: $\{\alpha,-\alpha\}$), revealing DRAG-blocked frequencies at $\pm\alpha$. The green line corresponds to the pulse with dual-DRAG plus an additional third DRAG at $\alpha$ (DRAG set: $\{\alpha,-\alpha,\alpha\}$). The red line represents the pulse with two DRAGs at the same detuning $\alpha$ (DRAG set: $\{\alpha,\alpha\}$), resulting in a larger frequency shift of approximately 170 MHz, compared to the smaller shift of about 50 MHz seen in the green line.
In summary, the result from our Eq. (35) is derived within the DRAG framework using a full quantum mechanical analysis. This allowed us to obtain both the correction formula and the effective Hamiltonian, which together explain the high performance of HD DRAG. Our findings are consistent with the spectrum optimization approach presented in reference [17].
@@ VI Fitting function for the excitation rate
$\begin{split}\bar{C}_{i}&=\frac{1}{2\pi}\int_{\phi=0}^{2\pi}e^{iH(\phi)\theta}d\phi,\\ H(\phi)&=\cos{\phi}\sigma_{x}+\sin{\phi}\sigma_{y},\end{split}$ (43)
@@
where $p_{1}(m-1)$ is the population on the excited state after m1m-1 Clifford gates. We can get the population of the excited state in $\rho_{m}$ as
@@ VII Explanation of the inter-pulse interference effects
The ExPC as a function of gate time $t_{g}$ shows an oscillatory behavior in Fig. 4(c) of the main text. This phenomenon can be explained by the inter-pulse interference effect. As shown in Fig. 8, the red line represents twice the simulated excitation rate of the $\rm\sqrt{X}$ gate over gate time, the orange line shows the results from a single U3 gate, and the blue line corresponds to the result from three U3 gates. As the pulse duration increases, the decreasing trend of ExPC for three U3 gates is not monotonic, and oscillations appear, which are attributed to the interference effects between the different pulses. In an RB circuit, which consists of hundreds of U3 gates, the fitted excitation rate is influenced by these inter-pulse interference effects. The practical excitation rate lies within the oscillation range, as shown by the comparison among the simulated ExPCs for $\rm\sqrt{X}$, $\rm U3$, and $\rm U3\times 3$ in Fig. 8.
@@ VIII Comparison of different DRAG sets
To compare the suppression of spurious transitions with different DRAG sets, we present the experimental results for RB and Ramsey interference error filter circuit with four different DRAG sets, as shown in Fig. 9. The ExPC is significantly reduced with the recursive DRAG scheme, and the best suppression occurs with the optimized dual-DRAG, using the DRAG set $\{\alpha,\Delta_{\rm opt},-\Delta_{\rm opt}\}$, as shown in Fig. 9(b). The significant suppression is further confirmed in Fig. 9(c). The irregular lines in the top panel may be attributed to noise in the system.
@@ X Suppression of the spurious transitions
XI Suppression of the spurious transitions with multiple dual-DRAGs
Figure 12: The target qubit $\rm Q_{0}$ couples to two spectator qubits (or TLSs) as shown in the top of the figure, with coupling strengths $g_{01}/2\pi=0.8\,\mathrm{MHz}$ between $\rm Q_{0}$ and $\rm Q_{1}$ and $g_{02}/2\pi=2.5\,\mathrm{MHz}$ between $\rm Q_{0}$ and $\rm Q_{2}$. The normalized time-domain driving pulses are shown in panels (a, e, i), where blue lines represent pulses without dual-DRAG corrections and green, red, and orange lines show pulses with dual-DRAG corrections, respectively. The corresponding frequency-domain representations are presented in panels (b, f, j). Simulation results for the pulse sequence from Fig. 3(b) (main text) with N=20 are displayed in panels (c, d, g, h, k, l), where the spurious transitions to spectator qubits (or TLSs) are amplified - results for $\rm Q_{1}$ are shown in (c, g, k) while those for $\rm Q_{2}$ are in (d, h, l). The detunings are defined as $\Delta_{1}=\omega_{1}-\omega_{0}$ and $\Delta_{2}=\omega_{2}-\omega_{0}$, with specific values: for panels (a-d), $\Delta_{1}/2\pi=40\,\mathrm{MHz}$ and $\Delta_{2}/2\pi=90\,\mathrm{MHz}$; for (e-h), $\Delta_{1}/2\pi=80\,\mathrm{MHz}$ and $\Delta_{2}/2\pi=90\,\mathrm{MHz}$; and for (i-l), $\Delta_{1}/2\pi=40\,\mathrm{MHz}$ and $\Delta_{2}/2\pi=50\,\mathrm{MHz}$.
The dual-DRAG protocol can be generalized to suppress multiple unwanted transitions, though several important considerations must be noted. Multiple DRAG corrections can be applied recursively at different detuning frequencies to eliminate corresponding spectral components of the control pulse. However, standard DRAG pulses may become difficult to calibrate when addressing small detunings due to strong spectral distortions. Our dual-DRAG approach, employing two mirrored DRAG operations for each target frequency, improves this situation and expands the applicable parameter range. This extension comes at the cost of requiring twice as many DRAG operations, with the overall performance strongly dependent on the chosen detuning frequencies.
To illustrate this capability, we examine a system where the target qubit $\rm Q_{0}$ couples to two spectator qubits $\rm Q_{1}$ and $\rm Q_{2}$ with detunings $\Delta_{1}=\omega_{1}-\omega_{0}$ and $\Delta_{2}=\omega_{2}-\omega_{0}$, where $\omega_{i}$ denotes the frequency of ${\rm Q}_{i}$ ($i=0,1,2$), as depicted on the top of figure 12.
Case 1: Two spectators with detunings $\Delta_{1}/2\pi=40\,\mathrm{MHz}$ (small) and $\Delta_{2}/2\pi=90\,\mathrm{MHz}$ (large). For a gate time $t_{g}=25\,\mathrm{ns}$, we apply dual-DRAG corrections at $\pm 40\,\mathrm{MHz}$ and $\pm 90\,\mathrm{MHz}$, followed by fine-tuning. Figure 12(a) compares the pulse shapes without and with dual-DRAG in the time domain, while Fig. 12(b) shows their frequency-domain counterparts. As demonstrated in Figs. 12(c, d), the dual-DRAG corrections suppress spurious transitions to both $\rm Q_{1}$ and $\rm Q_{2}$ while maintaining a low gate error with $\epsilon=4\times 10^{-5}$.
Case 2: Near-resonant spectators $\Delta_{1}/2\pi=80\,\mathrm{MHz}$, $\Delta_{2}/2\pi=90\,\mathrm{MHz}$. For closely spaced spectators, the dual-DRAG scheme remains effective. The pulse shapes (time and frequency domains) are shown in Figs. 12(e) and (f), respectively. Figures 12(g, h) confirm significant suppression of spurious transitions to $\rm Q_{1}$ and $\rm Q_{2}$, with a gate error of $\epsilon=1\times 10^{-5}$.
Case 3: Two small detunings $\Delta_{1}/2\pi=40\,\mathrm{MHz}$, $\Delta_{2}/2\pi=50\,\mathrm{MHz}$. For smaller detunings, spurious transitions are still suppressed, but the driving pulse exhibits stronger frequency-domain distortion, where the amplitude of the side lobes around 4.9 GHz and 5.1 GHz exceeds that of the main lobe (Fig. 12(j)). While the gate error increases to $4\times 10^{-3}$, prolonging the gate time can mitigate pulse distortion and restore fidelity.
The performance of multiple dual-DRAG operations is primarily determined by the detuning frequencies involved. While applying multiple dual-DRAG corrections at large detunings (typically $\Delta/2\pi\gtrsim 50$  MHz for 25-ns pulses) remains straightforward, small-detuning cases ($\Delta/2\pi\lesssim 50$  MHz) present challenges. Although these can effectively suppress unwanted transitions, they may compromise gate fidelity due to increased spectral distortion. In such cases, gate performance can be improved by optimizing the gate duration.
XII Leakage suppression with recursive DRAG